The region in the first quadrant which is bounded by the curve y² =


4x, and the lines x = 4 and y = 0, is revolved about the x-axis.


Locate the centroid of the resulting solid of revolution.

1
Expert's answer
2022-06-09T14:02:10-0400

ANSWER The centroid is "(\frac{8}{3},0,0)"

EXPLANATION

If "x_{0}" is fixed, then when rotating the segment "\left \{\left ( x_{0},y,0 \right ) :0\leq y\leq 2\sqrt{x_{0}} \right \}" around x-axis, we get a circle : "y^{2}+z^{2}=4x_{0}" . So, the resulting solid of revolution is

"R=\left \{ \left ( x,y,z \right ): 0\leq x\leq 4,\, 0\leq y^{2} +z^{2}\leq 4x\right \}"



The centroid "(x_{C},y_{C},z_{C})" is given by

"x_{C}=\frac{\iiint_{R}xdV} {\iiint_{R}dV} ,\, \,y_{C}=\frac{\iiint_{R}ydV} {\iiint_{R}dV} ,\, \,z_{C}=\frac{\iiint_{R}zdV} {\iiint_{R}dV}"

"\iiint_{R}dV=\int_{0}^{4}dx\iint_{S_{x}} dydz" ,

where "S_{x}" is a circle whose radius is "4x" :

"S_{x }= \left \{ \left ( y,z \right ) :y^{2}+z^{2}\leq 4x\right \}"

"Area \, \, of\, \, S_{x}=\iint_{S_{x}} dydz=4\pi\cdot x",

so "\iiint_{R}dV=4\pi\int_{0}^{4}xdx=4\pi\left [ \frac{x^{2}}{2} \right ]_{0}^{4}=2\pi\left (16-0 \right )=32\pi"

"\iiint_{R}xdV=\int_{0}^{4}xdx\iint_{S_{x}} dydz=4\pi\int_{0}^{4} x ^{2}=4\pi\left [ \frac{x^{3}}{3} \right ]_{0}^{4}=\frac{256\pi}{3} \Rightarrow x_{C}=\frac{8}{3} \cong 2.6667"

Now, to calculate integrals over the region "S_{x}" we replace the Cartesian coordinates with polar coordinates : "y=r\cos t, z=r\sin t" "(0\leq t\leq 2\pi, 0\leq r\leq 2\sqrt{x})" . Therefore

"\iiint_{R}ydV=\int_{0}^{4} dx\iint_{S_{x}}y dydz= \int_{0}^{4} \left [ \int_{0}^{2\pi} \int_{0}^{2\sqrt{x}}r^{2}\cos tdrdt\right ]dx=\\=\left ( \int_{0}^{2\pi}\cos tdt \right )\left ( \int_{0}^{4}\int_{0}^{2\sqrt{x}}r^{2}drdx \right )=0 ,"

because "\int_{0}^{2\pi}\cos tdt =0."

"\iiint_{R}zdV=\int_{0}^{4} dx\iint_{S_{x}}z dydz= \int_{0}^{4} \left [ \int_{0}^{2\pi} \int_{0}^{2\sqrt{x}}r^{2}\\ \sin tdrdt\right ]dx=\\=\left ( \int_{0}^{2\pi}\\ \sin dt \right )\left ( \int_{0}^{4}\int_{0}^{2\sqrt{x}}r^{2}drdx \right )=0 ,"

since "\int_{0}^{2\pi} \sin tdt =0" . Hence "y_{C}=0, z_{C}=0"